# Pattern formation and spatiotemporal chaos in relativistic degenerate plasmas

### v1

@@ III Conditions for the existence of spatiotemporal chaos
A perturbation in the form of a wave $A(x,t)$ is said to exhibit spatiotemporal chaos [13, 15] if the following conditions are satisfied.
@@
. We study the temporal dynamics of $A(x,t)$ and calculate the largest Lyapunov exponent spectra of the time series of $A(x,t)$ for all real values of $x$ by using the method developed by Rosenstein [11]. The positive values of the largest Lyapunov exponent in the entire domain of $x$ confirm the existence of temporal chaos in the system of Eqs. (4) and (5).
@@
Typically, the statistical independence is estimated by the decay of two-point correlation function $C(r)$ or the decay of the mutual information function $I(r)$ with increasing values of the distance $r$ between two points. While the former is only a necessary condition, the vanishing of $I(r)$ is both a necessary and sufficient condition for the statistical independence. We first define the correlation function as [12, 13, 9]
@@
[…]
@@
Next, we define the mutual information function as [14, 15]
[…]
where the distributions $p_{x_{1},x_{2}}(u_{1},u_{2})$, $p_{x_{1}}(u_{1})$, and $p_{x_{2}}(u_{2})$ are generated by the time series of $\{A(x_{1},t)\}$ and $\{A(x_{2},t)\}$ for all $t$. Intuitively, $p_{x_{1},x_{2}}(A(x_{1}),A(x_{2}))dA(x_{1})dA(x_{2})$ is the fraction of time that both $A(x_{1},\cdot)\in(A(x_{1}),A(x_{1})+dA(x_{1}))$ and $A(x_{2},\cdot)\in(A(x_{2}),A(x_{2})+dA(x_{2}))$, and $p_{x_{1}}(A)dA$ is the fraction of time that $A(x_{1},\cdot)\in(A,A+dA)$, etc. We employ the Gaussian joint probability density function to calculate the mutual information for the present system of Eqs. (4) and (5).
IV Numerical simulation: Evolution of spatiotemporal chaos
@@
From Eq. (11) and the discussion in Sec. II, we note that the modulation perturbation is unstable in $0<k<k_{c}$ and stable for $k>k_{c}$ such that $k=k_{c}$ defines the pitchfork bifurcation curve. The instability growth rate tends to increase in $0<k<k_{m}$ (achieving a maximum value at $k=k_{m}$) and decreases in $k_{m}<k<k_{c}$ (having a cut-off at $k=k_{c}$). In this regime of small growth rate, the EM waves are essentially trapped by the density humps (similar to Fig. 2), and a stationary solution of Eqs. (4) and (5) in a frame $\xi=x-v_{g}t$ moving with the group velocity $v_{g}$ can be obtained as $N=3b_{3}|A|^{2}/(1-v_{g}^{2})\sim\rm{sech}^{2}\left({\xi/w}\right)$ for some constant $w$. Clearly, $v_{g}<1$ (or the group velocity is to be smaller than the sound speed) for a soliton solution to exist. Thus, below and sufficiently close to the bifurcation curve, i.e., in the region $k_{m}<k\lesssim k_{c}$, the dynamics of coupled EM waves and EAWs is truly subsonic in the sense that the MI growth rate is small $(\Gamma/k\ll 1)$ [10, 5]. So, a critical wave number $k_{c1}$ may exist in $k_{m}<k<k_{c}$ such that the motion of EM wave envelopes will be in temporal periodic or quasi-periodic states [8] if the modulation wave number lies in $k_{m}<k_{c1}<k<k_{c}$, and the coexistence of temporal chaos and coherent solitary patterns may occur if $k$ lies in $k_{m}<k<k_{c1}<k_{c}$ [8]. On the other hand, the region of $k$ below $k_{m}$ may not be subsonic if $k_{c}$ is relatively large, and Eq. (11) suggests that not only is the perturbed mode with the wave number $k$ unstable but also a higher-harmonic mode with the wave number $2k$. If $k$ is further reduced from $k_{m}$, there is a possibility of the existence of more and more higher harmonic modes that are unstable. Thus, the region $k_{m}<k<k_{c}$ may be appropriate for describing a few unstable modes [5]. We are, however, interested in the region $0<k<k_{m}$ where many solitary patterns may be excited through the modulational instability and saturated after collisions, fusions, and interactions with electron-acoustic density perturbations. In the numerical simulation, we excite the solitary patterns at the modulation wavelengths $l_{m}=L_{x}/m$, where $m=1$ is for the master mode and $m=2,3,\ldots,M$ are for the unstable harmonic modes due to pattern selection such that $M<[k^{-1}]$ [9].
@@
We solve Eqs. (4) and (5) numerically by considering a simulation box size $L_{x}=2\pi/k$ as the resonant wavelength for a perturbed mode, the number of grid points as $2048$, and the initial conditions as [5, 16, 9]
@@
where $A_{0}$ is the amplitude of the pump EM wave envelope and $b$ is a suitable constant, and we take $L\sim 500$ to mean relatively a small perturbation. Equations (4) and (5) are advanced in time using the Runge-Kutta scheme with a time step $dt\sim 10^{-4}$ or less such that $dt<dx^{2}/B_{0}$ is satisfied for the convergence of the scheme. We approximate the spatial derivatives by using the centered second-order difference formulas and assume $A=N=\partial^{2}A/\partial x^{2}=\partial^{2}N/\partial x^{2}=0$ at the boundaries [5]. Furthermore, we choose $A_{0}\sim 1$ and $b\sim 0.5$ and the fixed values as $K=0.01$ and $\alpha=0.03$. Thus, the key parameters are the modulation wave number $k$ and the degeneracy parameter $R_{0}$. While the former is responsible for the excitation of different (master and harmonic) modes in the pattern selection, the latter corresponds to the regimes of degenerate dense plasmas. Typically, the degeneracy effect starts playing a role for $R_{0}\gtrsim 1$, i.e., in the regimes of electron number density close to or exceeding the critical number density $n_{\rm{cr}}=6\times 10^{29}~{}\rm{cm}^{-3}$. Thus, values of $R_{0}$ close to $5$, $25$, and $85$ or a bit more may, respectively, be roughly called moderate, strong, and ultra-relativistic regimes. In Secs. IV.1-IV.3, we will mainly consider these three regimes separately. Since the growth rate and the instability domain are seen to be significantly reduced for $R_{0}<1$ (See the dash-dotted line of Fig. 1), the case of weak relativistic degeneracy is not of interest and thus not discussed in the present study.
IV.1 Moderate relativistic degeneracy
@@
We consider a relativistic degeneracy regime with $R_{0}\sim 5$ for degenerate electrons that corresponds to the density: $n_{\rm{d0}}\sim 7.5\times 10^{31}~{}\rm{cm}^{-3}$. Figure 2 shows the spatial profiles of the perturbed EM wave field $|A|$ (solid line) and the number density perturbation of degenerate electrons $N$ associated with the EAWs (dash-dotted line) for $R_{0}=5$ and different values of the modulation wave number: $k=0.15$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.043$ [subplot (c)] within the domain $0<k<k_{m}$ after some small interval of times, $t=12,~{}14$, and $30$ respectively, to observe the initial excitation of master and harmonic modes. We observe that the amplitudes of perturbations grow beyond the unity due to the modulational instability, and an excited EM field $(|A|)$ is highly correlated with the density fluctuation $(N)$ having the highest peaks near the center at $x=0$. Furthermore, in the moderate degeneracy regime, the EM waves are trapped by the density humps. Subplot (a) of Fig. 2 shows that the pattern selection with $k=0.15$ leads to the excitations of three ($<[k^{-1}]=6$) modes: one master mode with wavelength $l_{m}=L_{x}$, initially peaked at $x=0$ and two harmonic modes with wavelength $l_{m}=L_{x}/4$ initially peaked at two different points: $x=-L_{x}/4$ and $L_{x}/4$ (i.e., near the points $x=-10$ and $x=10$ respectively) in the simulation box $[-L_{x}/2,~{}L_{x}/2]$. It is seen that the master pattern is first formed (near $x=0$) and the harmonic modes begin to appear near $x=\pm 10$ from the master mode. As we successively reduce the modulation wave number from $k=0.15$ to $k=0.11$ and $k=0.043$, the excitation of more harmonic modes (four and thirteen) with reduced wavelengths $l_{m}=L_{x}/6$ [See subplot (b)] and $l_{m}=L_{x}/14$ [See subplot (c)] are seen. The density fluctuations with $N<0$ also appear in Fig. 2, which indicates that during the pattern formation, electrons are pushed back by the EM wave-driven ponderomotive force from the regions of the master mode and harmonic waves. Such density-depleted electrons move stochastically on either side of the patterns, and thus cannot arrest the EM wave fields [5].
@@
As time goes on (i.e., going beyond the times $t=12,~{}14$, and $30$ for initial excitation of modes), some new interesting phenomena are observed. The EM master mode and harmonic modes excited at different points of $x$ with large scales collide themselves to generate higher harmonic modes with short wavelengths and also fuse after interactions due to strong EAW emission. The contour plots (Fig. 3) corresponding to Fig. 2, show that for $k=0.15$ [subplot (a)] the EM master pattern (excited near $x=0$) and the two harmonic patterns (excited near $x=\pm 10$) move stochastically, i.e., their amplitudes oscillate and widths vary in time, and after some time, i.e., at $t\approx 250$, the harmonic modes collide themselves and eventually fuse to generate a new incoherent pattern due to electron-acoustic radiation. The system energy is so weak that the new pattern does not propagate any more and eventually it disappears after $t\approx 300$. The system is then said to be in TC state as evident from the positive values of the largest Lyapunov exponent shown in the subplot (a) of Fig. 4. However, the spatial behaviors of the patterns are still in the SPC state due to EAW emission as evident from Fig. 5 that the correlation function $C(r)$ [subplot (a); see the solid line] still lies in the interval $-0.1<C(r)<0.1$ and the mutual information $I(r)$ [subplot (b); see the solid line] does not approach to zero as $r$ gradually increases. It follows that a few patterns may not be sufficient to cause STC, i.e., the coherence of EM waves is still partially retained and the system is in the coexistence of TC and SPC states. Subplot (b) of Fig. 3 shows that as $k$ is further reduced from $k=0.15$ to $k=0.11$, the pattern selection leads to the excitations of five modes: four harmonic modes at $x=\pm L_{x}/3,~{}\pm L_{x}/6$ and a master mode at $x=0$ with initial wavelengths $L_{x}/6$ and $L_{x}$ respectively. It is observed that the patterns excited at $x=\pm L_{x}/3$ disappear after $t\approx 40$ due to weak EM wave energy and the influence of the strong electron-acoustic radiation. However, the other two harmonic modes initially excited at $x=\pm L_{x}/6$ collide at two different times $t\approx 200$ and $t\approx 580$ with the master mode (initially excited near $x=0$) and eventually fuse to generate a single incoherent distorted master mode with a higher amplitude but a narrower width. From subplot (b) of Fig. 4, the largest Lyapunov exponent is seen to be positive. Also, from subplots (a) and (b) of Fig. 5 (See the dash-dotted lines) we find an exponential decay of both the correlation function $[C(r)]$ and the mutual information $[I(r)]$. The correlation and the mutual information decay lengths, respectively, are $\xi_{C}\approx 9.2336$ and $\xi_{I}\approx 2.6781$ (See Table 1), i.e., $\xi_{C}$, $\xi_{I}<<2048$, the system grid size. So, the system is said to emerge in STC. Thus, even in the moderate relativistic regime, a few number initially excited modes can lead to the emergence of STC due to strong electron-acoustic radiation.
@@
Figure 2: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation ($N$) are shown over the spatial domain in the moderate relativistic degenerate regime with $R_{0}=5$. It is seen that the pattern selections lead to the excitations of three, five and thirteen solitary modes corresponding to $k=0.15$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.043$ [subplot (c)] respectively.
@@
We also consider the case, when many unstable solitary patterns can be excited and saturated in the pattern selection, e.g., for $k=0.043$ and to verify the emergence of STC in the moderate degenerate regime. Figure 3 (c) shows that thirteen solitary patterns can be initially formed at points $x=\pm L_{x}/14,~{}\pm l_{x}/7,~{}\pm 3L_{x}/14,~{}\pm 2L_{x}/7,~{}5L_{x}/14,~{}\pm 3L_{x}/7$, and $x=0$ due to the pattern selection for $k=0.043$. The first collision occurs around $t=90$ between the modes excited at $x=3L_{x}/14$ and $145L_{x}/$ 14, and thereafter two binary collisions occur at $t\approx 190$. Collisions also take place after $t=200$ in which new incoherent patterns are formed accompanied by strong electron-acoustic radiation. It is observed that three harmonic modes initially excited at $x=-3L_{x}/14$, $-2L_{x}/7$, and $-5L_{x}/14$ collide and fuse to generate a new higher harmonic mode with a higher amplitude but narrower width. Also, two other modes initially excited at $x=-3L_{x}/7$ and $72L_{x}/$ 7 disappear shortly due to influence of the electron-acoustic radiation. As time goes on, repeated collisions among the harmonic modes occur and the new incoherent patterns get distorted. We note that the harmonic mode initially excited at $x=-L_{x}/14$ collides with the master mode and again collides with the new incoherent pattern formed by the repeated collisions of four solitary patterns initially excited at $x=L_{x}/14,~{}L_{x}/7,~{}3L_{x}/14$, and $-5L_{x}/14$. Finally, the thirteen solitary patterns with long wavelengths fuse into four incoherent patterns with small wavelengths. Form subplot (c) of Fig. 4, we note that the largest Lyapunov exponent is positive and from subplots (a) and (b) of Fig. 5 (dotted lines), both the correlation function and the mutual information tend to assume zero values with increasing values of $r$. The correlation and mutual information decay lengths, respectively, are $\xi_{C}\approx 4.1894$ and $\xi_{I}\approx 1.4613$ (See Table 1), which are well below the system grid size $2048$. Thus, it may be inferred that the state of STC of the system emerges. In the STC state, the system energy, which was initially stored in thirteen solitary modes, is now spatially redistributed to four higher harmonic incoherent modes with short wavelengths in the process of random collisions and fusions among the patterns. Here, we note that a critical value $k_{\rm{cs}}$ of $k$ must exist below which a transition from TC to STC can occur. In the case of $R_{0}=5$, such a value lies in $0.11\lesssim k_{cs}<0.15<k_{m}$. Thus, in the regime of moderate relativistic degeneracy of electrons, it may be concluded that within the instability domain (where the growth rate tends to increase with $k$) but below the wave number of modulation (perturbation) $k_{m}$ at which the instability growth rate is maximum, the nonlinear interaction of circularly polarized EM waves and electron-acoustic density perturbations can no longer give rise to coherent solitons but lead to the emergence of spatiotemporal chaos in which energy transfer indeed occurs from initially excited many coherent harmonic modes to a few new higher harmonic incoherent modes with short wavelengths.
IV.2 Strong relativistic degeneracy

### v2

@@ 3 Conditions for the existence of spatiotemporal chaos
A perturbation in the form of a wave $A(x,t)$ is said to exhibit spatiotemporal chaos [cai2000chaotic, cai2001spatiotemporal] if the following conditions are satisfied.
@@
We study the temporal dynamics of $A(x,t)$ and calculate the largest Lyapunov exponent spectra of the time series of $A(x,t)$ for all real values of $x$ by using the method developed by Rosenstein [rosenstein1993practical]. The positive values of the largest Lyapunov exponent in the entire domain of $x$ confirm the existence of temporal chaos in the system of Eqs. (4) and (5).
@@
Typically, the statistical independence is estimated by the decay of two-point correlation function $C(r)$ or the decay of the mutual information function $I(r)$ with increasing values of the distance $r$ between two points. While the former is only a necessary condition, the vanishing of $I(r)$ is both a necessary and sufficient condition for the statistical independence. We first define the correlation function as [hohenberg1989chaotic, cai2000chaotic, he2002harmonic]
@@
where the angular brackets $\left$$\cdot\right$> denote the temporal mean. If $C(r)$ behaves as $\exp(-r/\xi)$ for $r\to 0$ and $\xi\ll L_{x}$, or $C(r)\to 0$ for some $\xi$ and $r$, where $\xi$ is the correlation length and $L_{x}$ is the system size, then the system is said to be statistically independent or spatially incoherent. However, as in Ref. [he2002harmonic], if the correlation function lies in $0<C(r)<1$, the system may be in the spatially partial coherent state. In general, $C(r)$ lies within 1<C(r)<1-1<C(r)<1 but may exponentially decay to zero for a particular system.
@@
Next, we define the mutual information function as [blahut1987principles, cai2001spatiotemporal]
$\begin{split}&I(r)\equiv I(x_{1},x_{2})\\ &=\int p_{x_{1},x_{2}}(u_{1},u_{2})\log\left[\frac{p_{x_{1},x_{2}}(u_{1},u_{2})}{p_{x_{1}}(u_{1})p_{x_{2}}(u_{2})}\right]du_{1}du_{2},\end{split}$ (15)
where the distributions $p_{x_{1},x_{2}}(u_{1},u_{2})$, $p_{x_{1}}(u_{1})$, and $p_{x_{2}}(u_{2})$ are generated by the time series of $\{A(x_{1},t)\}$ and $\{A(x_{2},t)\}$ for all $t$. Intuitively, $p_{x_{1},x_{2}}(A(x_{1}),A(x_{2}))dA(x_{1})dA(x_{2})$ is the fraction of time that both $A(x_{1},\cdot)\in(A(x_{1}),A(x_{1})+dA(x_{1}))$ and $A(x_{2},\cdot)\in(A(x_{2}),A(x_{2})+dA(x_{2}))$, and $p_{x_{1}}(A)dA$ is the fraction of time that $A(x_{1},\cdot)\in(A,A+dA)$, etc. We employ the Gaussian joint probability density function to calculate the mutual information for the present system of Eqs. (4) and (5). We note that while Lyapunov exponents typically quantify the exponential divergence of nearby trajectories in a chaotic system, with positive values indicating chaos and short-term predictability, correlation functions measure the relationship between a system’s past and present states, exhibiting long-range order in chaotic systems and their complexity. On the other hand, mutual information assesses the statistical dependency between two variables, helping to understand information flow and complexity within a chaotic system.
4 Numerical simulation: Evolution of spatiotemporal chaos
@@
From Eq. (11) and the discussion in Sec. 2, we note that the modulation perturbation is unstable in $0<k<k_{c}$ and stable for $k>k_{c}$ such that $k=k_{c}$ defines the pitchfork bifurcation curve. The instability growth rate tends to increase in $0<k<k_{m}$ (achieving a maximum value at $k=k_{m}$) and decreases in $k_{m}<k<k_{c}$ (having a cut-off at $k=k_{c}$). In this regime of small growth rate, the EM waves are essentially trapped by the density humps (similar to Fig. 2), and a stationary solution of Eqs. (4) and (5) in a frame $\xi=x-v_{g}t$ moving with the group velocity $v_{g}$ can be obtained as $N=3b_{3}|A|^{2}/(1-v_{g}^{2})\sim\rm{sech}^{2}\left({\xi/w}\right)$ for some constant $w$. Clearly, $v_{g}<1$ (or the group velocity is to be smaller than the sound speed) for a soliton solution to exist. Thus, below and sufficiently close to the bifurcation curve, i.e., in the region $k_{m}<k\lesssim k_{c}$, the dynamics of coupled EM waves and EAWs is truly subsonic in the sense that the MI growth rate is small $(\Gamma/k\ll 1)$ [rizzato1998solitons, banerjee2010spatiotemporal]. So, a critical wave number $k_{c1}$ may exist in $k_{m}<k<k_{c}$ such that the motion of EM wave envelopes will be in temporal periodic or quasi-periodic states [he2000pattern] if the modulation wave number lies in $k_{m}<k_{c1}<k<k_{c}$, and the coexistence of temporal chaos and coherent solitary patterns may occur if $k$ lies in $k_{m}<k<k_{c1}<k_{c}$ [he2000pattern]. On the other hand, the region of $k$ below $k_{m}$ may not be subsonic if $k_{c}$ is relatively large, and Eq. (11) suggests that not only is the perturbed mode with the wave number $k$ unstable but also a higher-harmonic mode with the wave number $2k$. If $k$ is further reduced from $k_{m}$, there is a possibility of the existence of more and more higher harmonic modes that are unstable. Thus, the region $k_{m}<k<k_{c}$ may be appropriate for describing a few unstable modes [banerjee2010spatiotemporal]. We are, however, interested in the region $0<k<k_{m}$ where many solitary patterns may be excited through the modulational instability and saturated after collisions, fusions, and interactions with electron-acoustic density perturbations. In the numerical simulation, we excite the solitary patterns at the modulation wavelengths $l_{m}=L_{x}/m$, where $m=1$ is for the master mode and $m=2,3,\ldots,M$ are for the unstable harmonic modes due to pattern selection such that $M<[k^{-1}]$ [he2002harmonic].
@@
We solve Eqs. (4) and (5) numerically by considering a simulation box size $L_{x}=2\pi/k$ as the resonant wavelength for a perturbed mode, the number of grid points as $2048$, and the initial conditions as [banerjee2010spatiotemporal, misra2009pattern, he2002harmonic]
@@
where $A_{0}$ is the amplitude of the pump EM wave envelope and $b$ is a suitable constant, and we take $L\sim 500$ to mean relatively a small perturbation. Equations (4) and (5) are advanced in time using the Runge-Kutta scheme with a time step $\Delta t\sim 10^{-4}$ or less such that $\Delta t<\Delta x^{2}\min\left\{1/B_{0},1/2,1/6b_{3}\right\}$ (See Appendix A) is satisfied for the convergence of the scheme. We approximate the spatial derivatives by using the centered second-order difference formulas and assume periodic boundary condition as described in appendix A. Furthermore, we choose $A_{0}\sim 1$ and $b\sim 0.5$ and the fixed values as $K=0.01$ and $\alpha=0.03$. Thus, the key parameters are the modulation wave number $k$ and the degeneracy parameter $R_{0}$. While the former is responsible for the excitation of different (master and harmonic) modes in the pattern selection, the latter corresponds to the regimes of degenerate dense plasmas. Typically, the degeneracy effect starts playing a role for $R_{0}\gtrsim 1$, i.e., in the regimes of electron number density close to or exceeding the critical number density $n_{\rm{cr}}=6\times 10^{29}~\rm{cm}^{-3}$. Thus, values of $R_{0}$ close to $5$, $25$, and $85$ or a bit more may, respectively, be roughly called moderate, strong, and ultra-relativistic regimes. In Secs. 4.1-4.3, we will mainly consider these three regimes separately. Since the growth rate and the instability domain are seen to be significantly reduced for $R_{0}<1$ (See the dash-dotted line of Fig. 1), the case of weak relativistic degeneracy is not of interest and thus not discussed in the present study. We note that the MI domain for the evolution of EM wave envelopes is $0<k<k_{c}$ of which the subdomain $0<k\lesssim k_{m}$ corresponds to the region where the MI growth rate tends to increase and becomes high at $k_{m}$, and the excitations of large-scale EM solitary waves are possible. On the other hand, the region, $k_{m}<k<k_{c}$ corresponds to the solitonic region where the MI growth rate reduces with $k$ and gets saturated at $k=k_{c}$. This region exhibits the excitation of small-scale stable EM solitary waves. So, we are interested in the domain of $0<k\lesssim k_{m}$. In the pattern selection, we choose three different values of $k~(<k_{m})$, and we start with a value of $k$ close to but slightly lower than $k_{m}$, and successively reduce to choose two other values for the excitation of many solitary modes at large wavelengths. We will see that there appear mainly two subdomains of $0<k\lesssim k_{m}$: $0<k\lesssim k_{\rm{cs}}$ and $k_{\rm{cs}}<k\lesssim k_{m}$. While the former corresponds to the region for the emergence of STC, the latter is for the region where the system is in the coexistence of TC and SPC states. Here, $k_{\rm{cs}}$ is a critical value of $k~(<k_{m})$ at which a transition between these states occur. Furthermore, from Fig. 1, it is noted that the MI domain $0<k\lesssim k_{m}$ expands and $k_{m}<k<k_{c}$ contracts as the value of $R_{0}$ increases, i.e., the degeneracy regime shifts from a moderate to an ultra-relativistic regime, implying that the STC state is more likely to emerge in the ultra-relativistic degeneracy regimes with $R_{0}\gg 1$.
4.1 Moderate relativistic degeneracy
@@
We consider a relativistic degeneracy regime with $R_{0}\sim 5$ for degenerate electrons that corresponds to the density: $n_{\rm{d0}}\sim 7.5\times 10^{31}~\rm{cm}^{-3}$. In this case, the MI domain, where the growth rate tends to become high, is $0<k\lesssim k_{m}\equiv 0.18$ (cf. Fig. 1), and we choose three different values of $k~(<k_{m})$, namely $k=0.15$, $k=0.11$, and $k=0.043$ for the excitation of increasing number of wave modes ($k^{-1}]$) at different wavelengths due to the pattern selection. Figure 2 shows the spatial profiles of the perturbed EM wave field $|A|$ (solid line) and the number density perturbation of degenerate electrons $N$ associated with the EAWs (dash-dotted line) for $R_{0}=5$ and different values of the modulation wave number: $k=0.15$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.043$ [subplot (c)] within the domain $0<k<k_{m}$ after some small interval of times, t=12,14t=12,~14, and $30$ respectively, to observe the initial excitation of master and harmonic modes. We observe that the amplitudes of perturbations grow beyond the unity due to the modulational instability, and an excited EM field $(|A|)$ is highly correlated with the density fluctuation $(N)$ having the highest peaks near the center at $x=0$. Furthermore, in the moderate degeneracy regime, the EM waves are trapped by the density humps. Subplot (a) of Fig. 2 shows that the pattern selection with $k=0.15$ leads to the excitations of three ($k^{-1}]=6$) modes: one master mode with wavelength $l_{m}=L_{x}$, initially peaked at $x=0$ and two harmonic modes with wavelength $l_{m}=L_{x}/4$ initially peaked at two different points: $x=-L_{x}/4$ and $L_{x}/4$ (i.e., near the points x=10x=-10 and $x=10$ respectively) in the simulation box $L_{x}/2,~L_{x}/2]$. It is seen that the master pattern is first formed (near $x=0$) and the harmonic modes begin to appear near $x=\pm 10$ from the master mode. As we successively reduce the modulation wave number from $k=0.15$ to $k=0.11$ and $k=0.043$, the excitation of more harmonic modes (four and thirteen) with reduced wavelengths $l_{m}=L_{x}/6$ [See subplot (b)] and $l_{m}=L_{x}/14$ [See subplot (c)] are seen. The density fluctuations with $N<0$ also appear in Fig. 2, which indicates that during the pattern formation, electrons are pushed back by the EM wave-driven ponderomotive force from the regions of the master mode and harmonic waves. Such density-depleted electrons move stochastically on either side of the patterns, and thus cannot arrest the EM wave fields [banerjee2010spatiotemporal].
@@
As time goes on (i.e., going beyond the times t=12,14t=12,~14, and $30$ for initial excitation of modes), some new interesting phenomena are observed. The EM master mode and harmonic modes excited at different points of $x$ with large scales collide themselves to generate higher harmonic modes with short wavelengths and also fuse after interactions due to strong EAW emission. The contour plots (Fig. 3) corresponding to Fig. 2, show that for $k=0.15$ [subplot (a)] the EM master pattern (excited near $x=0$) and the two harmonic patterns (excited near $x=\pm 10$) move stochastically, i.e., their amplitudes oscillate and widths vary in time, and after some time, i.e., at $t\approx 250$, the harmonic modes collide themselves and eventually fuse to generate a new incoherent pattern due to electron-acoustic radiation. The system energy is so weak that the new pattern does not propagate any more and eventually it disappears after $t\approx 300$. The system is then said to be in TC state as evident from the positive values of the largest Lyapunov exponent shown in the subplot (a) of Fig. 4. However, the spatial behaviors of the patterns are still in the SPC state due to EAW emission as evident from Fig. 5 that the correlation function $C(r)$ [subplot (a); see the solid line] still lies in the interval 0.1<C(r)<0.1-0.1<C(r)<0.1 and the mutual information $I(r)$ [subplot (b); see the solid line] does not approach to zero as $r$ gradually increases. It follows that a few patterns may not be sufficient to cause STC, i.e., the coherence of EM waves is still partially retained and the system is in the coexistence of TC and SPC states. Subplot (b) of Fig. 3 shows that as $k$ is further reduced from $k=0.15$ to $k=0.11$, the pattern selection leads to the excitations of five modes: four harmonic modes at $x=\pm L_{x}/3,~\pm L_{x}/6$ and a master mode at $x=0$ with initial wavelengths $L_{x}/6$ and $L_{x}$ respectively. It is observed that the patterns excited at $x=\pm L_{x}/3$ disappear after $t\approx 40$ due to weak EM wave energy and the influence of the strong electron-acoustic radiation. However, the other two harmonic modes initially excited at $x=\pm L_{x}/6$ collide at two different times $t\approx 200$ and $t\approx 580$ with the master mode (initially excited near $x=0$) and eventually fuse to generate a single incoherent distorted master mode with a higher amplitude but a narrower width. From subplot (b) of Fig. 4, the largest Lyapunov exponent is seen to be positive. Also, from subplots (a) and (b) of Fig. 5 (See the dash-dotted lines) we find an exponential decay of both the correlation function $[C(r)]$ and the mutual information $[I(r)]$. The correlation and the mutual information decay lengths, respectively, are $\xi_{C}\approx 9.2336$ and $\xi_{I}\approx 2.6781$ (See Table 2), i.e., $\xi_{C}$, $\xi_{I}<<2048$, the system grid size. So, the system is said to emerge in STC. Thus, even in the moderate relativistic regime, a few number initially excited modes can lead to the emergence of STC due to strong electron-acoustic radiation.
@@
Figure 2: Initial excitations of master and harmonic solitary patterns for the EM wave field ($|A|$) and the associated density perturbation (NN) are shown over the spatial domain in the moderate relativistic degenerate regime with $R_{0}=5$. It is seen that the pattern selections lead to the excitations of three, five and thirteen solitary modes corresponding to $k=0.15$ [subplot (a)], $k=0.11$ [subplot (b)], and $k=0.043$ [subplot (c)] respectively. The other fixed parameter values are $A_{0}=2$ and $b=0.5$.
@@
We also consider the case, when many unstable solitary patterns can be excited and saturated in the pattern selection, e.g., for $k=0.043$ and to verify the emergence of STC in the moderate degenerate regime. Figure 3 (c) shows that thirteen solitary patterns can be initially formed at points $x=\pm L_{x}/14,~\pm l_{x}/7,~\pm 3L_{x}/14,~\pm 2L_{x}/7,~5L_{x}/14,~\pm 3L_{x}/7$, and $x=0$ due to the pattern selection for $k=0.043$. The first collision occurs around $t=90$ between the modes excited at $x=3L_{x}/14$ and $5L_{x}/14$, and thereafter two binary collisions occur at $t\approx 190$. Collisions also take place after $t=200$ in which new incoherent patterns are formed accompanied by strong electron-acoustic radiation. It is observed that three harmonic modes initially excited at $x=-3L_{x}/14$, $2L_{x}/7$, and $5L_{x}/14$ collide and fuse to generate a new higher harmonic mode with a higher amplitude but narrower width. Also, two other modes initially excited at $x=-3L_{x}/7$ and $2L_{x}/7$ disappear shortly due to influence of the electron-acoustic radiation. As time goes on, repeated collisions among the harmonic modes occur and the new incoherent patterns get distorted. We note that the harmonic mode initially excited at $x=-L_{x}/14$ collides with the master mode and again collides with the new incoherent pattern formed by the repeated collisions of four solitary patterns initially excited at $x=L_{x}/14,~L_{x}/7,~3L_{x}/14$, and $5L_{x}/14$. Finally, the thirteen solitary patterns with long wavelengths fuse into four incoherent patterns with small wavelengths. Form subplot (c) of Fig. 4, we note that the largest Lyapunov exponent is positive and from subplots (a) and (b) of Fig. 5 (dotted lines), both the correlation function and the mutual information tend to assume zero values with increasing values of $r$. The correlation and mutual information decay lengths, respectively, are $\xi_{C}\approx 4.1894$ and $\xi_{I}\approx 1.4613$ (See Table 2), which are well below the system grid size $2048$. Thus, it may be inferred that the state of STC of the system emerges. In the STC state, the system energy, which was initially stored in thirteen solitary modes, is now spatially redistributed to four higher harmonic incoherent modes with short wavelengths in the process of random collisions and fusions among the patterns. Here, we note that a critical value $k_{\rm{cs}}$ of $k$ must exist below which a transition from TC to STC can occur. In the case of $R_{0}=5$, such a value lies in $0.11\lesssim k_{cs}<0.15<k_{m}$. Thus, in the regime of moderate relativistic degeneracy of electrons, it may be concluded that within the instability domain (where the growth rate tends to increase with $k$) but below the wave number of modulation (perturbation) $k_{m}$ at which the instability growth rate is maximum, the nonlinear interaction of circularly polarized EM waves and electron-acoustic density perturbations can no longer give rise to coherent solitons but lead to the emergence of spatiotemporal chaos in which energy transfer indeed occurs from initially excited many coherent harmonic modes to a few new higher harmonic incoherent modes with short wavelengths.
4.2 Strong relativistic degeneracy